what steps would I take to find the value of x in the following eqution?
[6x+5]/3-3/7x=-2

Answer by stanbon(75887)

[6x+5]/3-3/7x=-2
-----------------
Multiply thru by 21x to get:
7[6x+5] - 3 = -42x
42x + 35 -3 = -42x
84x = -32
x = -8/21
===========================================
Cheers,
Stan H.